OpenStudy (anonymous):

Factor out the GCF from the polynomial 28x^7+16x^5-12x^3+20

OpenStudy (anonymous):

28=2*2*4 16=2*2*2*2 12=2*2*3 20=2*2*5 common to all terms is 2*2=4

OpenStudy (anonymous):

the answer would be 4?

OpenStudy (anonymous):

yes the answer is 4.